# Variations on cohomology rings and zero schemes

### v1

@@ 1 Introduction
The classical Poincaré–Hopf theorem allows one to relate Euler characteristic, a topological invariant of a manifold, to a local data extracted from the zeros of a vector field. As the seminal work of Carrell and Lieberman shows, one can extract more information if the manifold is a complex algebraic variety. Precisely, one has the following theorem [8]:
@@
In principle determining the filtration $F_{\bullet}$ is a difficult task. However, one can show [1, 2] that in certain situations it is induced by a grading, and hence $H^{*}(X,\mathbb{C})\simeq\mathbb{C}[Z]$. This is the case when the Borel subgroup of triangular matrices in $\operatorname{SL}_{2}$ acts on $X$, and the vector field $V$ is associated to the element $e=\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)\in\mathfrak{sl}_{2}$. After [5] we say in that case that the group acts regularly. By a result of Horrocks [17] if the vector field $V$ associated to an action of additive group has isolated zeros, it has in fact a unique zero. This seemingly strong condition holds for actions of principal $\operatorname{SL}_{2}$ on many well-studied spaces, including flag varieties, smooth Schubert varieties and Bott–Samelson resolutions.
@@
Assume that $\mathrm{B}_{2}$, the Borel subgroup of upper triangular matrices in $\operatorname{SL}_{2}$, acts regularly on a smooth projective variety $X$. Let $\mathcal{S}=e+\mathfrak{t}\subset\mathfrak{b}_{2}$. Let $\mathcal{Z}$ be the zero scheme of the vector field defined on $\mathcal{S}\times X$ by the infinitesimal action of the Lie algebra. Then $\mathcal{Z}\simeq\operatorname{Spec}H^{*}_{\mathbb{C}^{\times}}(X)$, where $H^{*}_{\mathbb{C}^{\times}}(X)$ is the cohomology ring of $X$ equivariant with respect to one-dimensional diagonal torus in $\mathrm{B}_{2}$. Moreover, the structure of algebra over $H^{*}_{\mathbb{C}^{\times}}(\operatorname{pt})\simeq\mathbb{C}[\mathfrak{t}]$ is given by the projection map $\mathcal{Z}\to\mathcal{S}\simeq\mathfrak{t}$. The grading on $H^{*}_{\mathbb{C}^{\times}}$ is determined by the weights of the $\mathbb{C}^{\times}$ representation on $\mathbb{C}[\mathcal{Z}]$. It comes from the action of $\mathbb{C}^{\times}$ on $\mathcal{Z}\subset\mathcal{S}\times X$ given by
@@
In [15] Hausel and the author extend the result to a much wider class of algebraic groups, called principally paired. The class includes all connected reductive groups and their parabolic subgroups. The affine plane $\mathcal{S}$ in general has to be replaced with a Kostant section inside the Lie algebra $\mathfrak{g}$ of the acting group $\mathrm{G}$. The torus action is analogous, and one gets the affine zero scheme $\mathcal{Z}$ over $\operatorname{Spec}H^{*}_{\mathrm{G}}(\operatorname{pt})$, and hence $\mathbb{C}[\mathcal{Z}]$ is a graded algebra over $H^{*}_{\mathrm{G}}(\operatorname{pt})$, proved to be canonically isomorphic to $H^{*}_{\mathrm{G}}(X)$. As before, the assumption is that $X$ is smooth projective, and the regular nilpotent has a unique zero. A version for singular varieties is also proved based on analogous result of [5]. The assumptions require the singular variety to be embedded in a smooth variety with regular action, and the restriction on cohomology rings has to be surjective.
@@
First, in Section 2.2, we show that the regularity assumption can be relaxed. We only ask for the zero scheme $\mathcal{Z}$ to be of the same dimension as the base $\mathcal{S}$, i.e equal to the rank of $\mathrm{G}$. We prove in Theorem 2.7 that the ring of functions on $\mathcal{Z}$ is still canonically isomorphic to the equivariant cohomology ring, as a graded algebra over $H^{*}_{\mathrm{G}}(\operatorname{pt})$. However, the cost of this generalisation is losing the affineness of $\mathcal{S}$, which only holds true if the action is regular. It then poses computational difficulties to potential calculations on cohomology rings. The examples are however plentiful, as we show that every spherical variety satisfies the conditions.
@@
In the same spirit, the results on equivariant cohomology in [15] where shown to hold true both for regular varieties and for GKM spaces, but without a clear connection between the two cases. However, we hope that the two results have a common generalisation into one common framework, both in cohomology and in K-theory case.
@@ 1.1 Notation
All the algebraic schemes in this paper, and in particular all the algebraic groups, are defined over complex numbers. Therefore for any scheme $X$, by $\mathbb{C}[X]=\Gamma(X,\mathcal{O}_{X})$ we denote the ring of global functions on $X$.
@@
For a linear algebraic group $\mathrm{G}$ with the Lie algebra $\mathfrak{g}$, we call an element $v\in\mathfrak{g}$ regular, if its centraliser $C_{\mathfrak{g}}(v)$ in $\mathfrak{g}$ has the dimension equal to the dimension of a maximal torus $\mathrm{T}$ in $\mathrm{G}$. If a regular element exists, the Lie algebra $\mathfrak{t}$ of $\mathrm{T}$ is a Cartan subalgebra, and its dimension is the rank of the group. The adjoint representation of $\mathrm{G}$ restricted to $\mathrm{T}$ splits into weight spaces. The weights are elements of $\mathfrak{t}^{*}$. The regular elements of $\mathfrak{t}$ are then those that are not annihilated by any of those weights. Recall that for $\mathrm{G}$ reductive those are called the roots of the group.
@@ 2.1 A short recap on Kostant sections and equivariant cohomology
Notice that the non-triviality of $\phi$ is simply equivalent to the non-triviality of $d\phi(h)$. However, a group can still be principally paired with $e$ mapping to zero. This is the case for any torus, where any closed one-parameter subgroup defines a morphism from $\mathrm{B}_{2}$, factoring via the quotient by the nilpotent part.
@@
We are given a homomorphism of $\mathfrak{b}_{2}=\operatorname{Lie}(\mathrm{B}_{2})$ to $\mathfrak{gl}_{n+1}$, given by
@@
The Kostant section defined in [15, 2.6.3] for a principally paired group $\mathrm{G}$ satisfies the following conditions.
@@
Whenever an algebraic group $\mathrm{G}$ acts on a smooth variety $X$, it yields an infinitesimal action of the Lie algebra $\mathfrak{g}$ of $X$. Formally, the action map $\mathrm{G}\times X\to X$ yields the map $T\mathrm{G}\times TX\to TX$ of the total spaces of the tangent bundles. Restricting to the $T_{1}\mathrm{G}\times X\simeq\mathfrak{g}\times X$ gives the Lie algebra action morphism $\mathfrak{g}\times X\to TX$.
Extending this by identity on $\mathfrak{g}$ yields $\mathfrak{g}\times X\to\mathfrak{g}\times TX$, which is a section of the vertical tangent bundle on $\mathfrak{g}\times X$, i.e. the tangent bundle $\mathfrak{g}\times TX$ in the direction of $X$. We call this vector field $V$ the total vector field of the action [^3] The author would like to thank Jakub Löwit for suggesting this slick definition. The definition in [15] is technically more involved.. For any $y\in\mathfrak{g}$ this restricts to a vector field on $X$, which we denote by $V_{y}$. We also consider the restriction $V_{\mathcal{S}}$ of $V$ to $\mathcal{S}\times X$.
@@
It is not known to the author whether $\mathcal{Z}_{\mathrm{tot}}$ itself has to be reduced under the conditions of the theorem. We touch this issue again in Sections 2.2 and 4.
@@
Now we need to see that the filtration is associated with the grading given by the weights of the $\mathbb{C}^{\times}$-action. First, for $t\in\mathbb{C}^{\times}$ let $t_{p}:\Omega^{p}_{X}\to\Omega^{p}_{X}$ be the pullback of forms along the map $t:X\to X$. We then construct the action of $\mathbb{C}^{\times}$ on the Koszul resolution via the following.
@@ 2.2 Zero schemes and equivariant cohomology for non-regular actions
The grading on $\mathbb{C}[\mathcal{Z}]$ is defined by the weight spaces of the $\mathbb{C}^{\times}$-representation thereon. That comes from the pullback of functions along the $\mathbb{C}^{\times}$ action on $\mathcal{Z}$ given by
@@
The assumption $\dim\mathcal{Z}=\dim\mathcal{S}$ implies that the generic fibres of the projection $\mathcal{Z}\to\mathcal{S}$ are finite, and hence a generic regular semisimple element has finitely many zeros. As the $\mathrm{T}$-fixed points are zeros of semisimple elements, there are only finitely many of them. If $X$ is a smooth projective variety with isolated torus fixed points, its cohomology is Tate. Hence we proved
@@
Here the localisation $c|_{\zeta_{i}}$ is an element of $H^{*}_{\mathrm{T}}(\operatorname{pt})=\mathbb{C}[\mathfrak{t}]$ and hence a function on $\mathfrak{t}$, which we apply to $w$. We then defined $\rho(c)$ as a function on the closed points of $\mathcal{Z}$, depending on a choice for each point.
@@
We already know the isomorphism $H^{*}_{\mathrm{T}}(X)\cong H^{0}(\mathcal{Z}_{\mathrm{B}}^{\prime},\mathcal{O}_{\mathcal{Z}_{\mathrm{B}}^{\prime}})$ and moreover $H^{*}_{\mathrm{G}}(X)=H^{*}_{\mathrm{T}}(X)^{\mathrm{W}}$. Hence to arrive at $H^{*}_{\mathrm{G}}(X)\simeq H^{0}(\mathcal{Z},\mathcal{O}_{\mathcal{Z}})$, we only have to prove that the Weyl action on $\mathcal{Z}_{\mathrm{B}}^{\prime}$ coming from the action on $\mathfrak{t}$ is identical with the topological action on $H^{*}_{\mathrm{T}}(X)$. This follows the proof of [15, Lemma 4.2]. By convention we let $\mathrm{W}$ act on the left on $\mathfrak{t}$, and on the right on $\mathbb{C}[\mathfrak{t}]$ and $H^{*}_{\mathrm{T}}(X)$.
@@
Note also that as for a torus $\mathrm{T}$ the regular (and only) nilpotent in $\mathfrak{t}=\operatorname{Lie}(\mathrm{T})$ is 0, no nontrivial variety is regular with $\mathrm{T}$-action. Therefore Theorem 2.4 cannot be applied directly.
One can notice that for a principally paired group $\mathrm{G}$ and a smooth projective $\mathrm{G}$-variety $X$, there are different levels of correspondence between the equivariant cohomology of $X$ and the ring of functions and coherent cohomology on $\mathcal{Z}$.
@@
It is not clear what should replace the question mark and a work is ongoing to determine under what assumptions this result holds. In fact, Löwit [21] proves a version of the claim for affine Bott–Samelson varieties which are not regular.
@@ 3 Singular varieties
For any $c\in H^{*}_{\mathrm{T}}(X)$ the function $\rho(c)$ is defined on $(e+v,p)$ by localisation to a torus-fixed point in the same $\mathrm{G}$-orbit as $p$. Hence the values of $\rho(c)$ on the points of $\mathcal{Z}^{Y}$ are defined by localisation only to torus-fixed points in $Y$. Therefore the restriction of $\rho(-)$ to $\mathcal{Z}^{Y}$ factors through $\iota^{*}$. Notice that in this step we are using the reducedness of $\mathcal{Z}^{Y}$, as this allows us to recover regular function from its values.
@@
A related work of Hausel [14] studies equivariant cohomology ring, as well as equivariant intersection cohomology of affine Schubert varieties. They are realised as rings of functions of zero schemes of commuting vector fields on a projective space. The definition of the vector fields is more involved.
@@ 4 K-theory conjecture for GKM spaces
In [15] a question was posed about extensions of Theorem 2.4 to other cohomology theories. The first non-trivial one, different from equivariant cohomology, is equivariant K-theory. Recall that the equivariant cohomology $H^{*}_{\mathrm{G}}(\operatorname{pt})$ of the point is equal to $\mathbb{C}[\mathfrak{t}]^{\mathrm{W}}=\mathbb{C}[\mathfrak{g}]^{\mathrm{G}}$. On the other hand, the equivariant algebraic K-theory $K^{0}_{\mathrm{G}}(\operatorname{pt})$ of the point is equal to the representation ring $R(\mathrm{G})$ of $\mathrm{G}$. After tensoring with $\mathbb{C}^{\times}$, this becomes isomorphic to $\mathbb{C}[\mathrm{T}]^{\mathrm{W}}=\mathbb{C}[\mathrm{G}]^{\mathrm{G}}$ [9, Theorem 6.1.4].
@@
Third, one could ask again how important the regularity assumption is. This is elaborated more in Remark 4.5 below. In this section, we prove a K-theory analogue of the GKM theorem for cohomology [15, Theorem 5.16].
@@
We can define $\rho$ by its values, as $\operatorname{Fix}^{\prime}_{\mathrm{T}}(X)$ is reduced. For any $\mathrm{T}$-linearised vector bundle $\mathcal{E}$ on $X$ let
@@
Note that if $0\to\mathcal{E}\to\mathcal{F}\to\mathcal{G}\to 0$ is an exact sequence of $\mathrm{T}$-linearised vector bundles, then for any $(t,p)\in\operatorname{Fix}^{\prime}_{\mathrm{T}}(X)$ we have $\operatorname{Tr}_{\mathcal{F}_{p}}(t)=\operatorname{Tr}_{\mathcal{E}_{p}}(t)+\operatorname{Tr}_{\mathcal{G}_{p}}(t)$. Therefore the function $\rho$ is well defined on $K^{0}_{\mathrm{T}}(X)$, and by definition it always gives a regular function on $\operatorname{Fix}_{\mathrm{T}}(X)$.
@@
We proved that $\tau$ is an injective map to $H\simeq K_{\mathrm{T}}(X)\otimes\mathbb{C}$, we only need to prove that $\tau\circ\rho=\operatorname{id}$. Take a class $\mathcal{E}]\in K^{0}_{\mathrm{T}}(X)$, where $\mathcal{E}$ is a $\mathrm{T}$-linearised vector bundle on $X$. Then $\tau\circ\rho([\mathcal{E}])$ is a function on $\mathrm{T}\times X^{\mathrm{T}}$ which at the point $(t,\zeta_{i})$ attains the value $\operatorname{Tr}_{\mathcal{E}_{\zeta_{i}}}(t)$. When we fix $\zeta_{i}$, this means that we restrict $\mathcal{E}$ to $\zeta_{i}$ to get a representation of $\mathrm{T}$, and then we compute the trace of $t$. But this is exactly how we define the isomorphism $K_{0}^{*}(\operatorname{pt})\cong\mathbb{C}[\mathrm{T}]$. Hence $\tau\circ\rho([\mathcal{E}])$, as an element on $H\subset\mathbb{C}[\mathrm{T}]^{s}$ is exactly the localisation of $\mathcal{E}$ to $X^{\mathrm{T}}$, which is the isomorphism between $K_{\mathrm{T}}(X)\otimes\mathbb{C}\cong H$. ∎
@@
We constructed the ring homomorphisms $\rho_{H}:H^{*}_{\mathrm{G}}(X)\to\mathbb{C}[\mathcal{Z}_{\mathrm{tot}}^{\operatorname{red}}]$ as in Theorem 2.4 and $\rho_{K}:K^{0}_{\mathrm{T}}(X)\otimes\mathbb{C}\to\mathbb{C}[\operatorname{Fix}^{\prime}_{\mathrm{T}}(X)]$ as in Theorem 4.2. Whenever they are both well-defined, the complex analytic map $\operatorname{ch}_{\mathrm{G}}$ (4.1) is compatible with the Chern character $\operatorname{ch}:K^{0}_{\mathrm{T}}(X)\otimes\mathbb{C}\to H^{*}_{\mathrm{G}}(X)$. This means that the following diagram commutes.
@@
and of the construction of Chern character from Chern roots.

### v2

@@ 1 Introduction
The classical Poincaré–Hopf theorem allows one to relate the Euler characteristic, a topological invariant of a manifold, to local data extracted from the zeros of a vector field. As the seminal work of Carrell and Lieberman shows, one can extract more information if the manifold is a complex algebraic variety. Precisely, one has the following theorem [8]:
@@
In principle determining the filtration $F_{\bullet}$ is a difficult task. However, one can show [1, 2] that in certain situations it is induced by a grading, and hence $H^{*}(X,\mathbb{C})\simeq\mathbb{C}[Z]$. This is the case when the Borel subgroup of triangular matrices in $\operatorname{SL}_{2}$ acts on $X$, and the vector field $V$ is associated to the element $e=\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)\in\mathfrak{sl}_{2}$. After [5] we say in that case that the group acts regularly. By a result of Horrocks [17], if the vector field $V$ associated to an action of the additive group has isolated zeros, it has in fact a unique zero. This seemingly strong condition holds for actions of principal $\operatorname{SL}_{2}$ subgroups on many well-studied spaces, including flag varieties, smooth Schubert varieties and Bott–Samelson resolutions.
@@
Assume that $\mathrm{B}_{2}$, the Borel subgroup of upper triangular matrices in $\operatorname{SL}_{2}$, acts regularly on a smooth projective variety $X$. Let $\mathcal{S}=e+\mathfrak{t}\subset\mathfrak{b}_{2}$, where $\mathfrak{t}$ is the diagonal Cartan subalgebra of $\mathfrak{sl}_{2}$. Let $\mathcal{Z}$ be the zero scheme of the vector field defined on $\mathcal{S}\times X$ by the infinitesimal action of the Lie algebra. Then $\mathcal{Z}\simeq\operatorname{Spec}H^{*}_{\mathbb{C}^{\times}}(X)$, where $H^{*}_{\mathbb{C}^{\times}}(X)$ is the cohomology ring of $X$ equivariant with respect to one-dimensional diagonal torus in $\mathrm{B}_{2}$. Moreover, the structure of algebra over $H^{*}_{\mathbb{C}^{\times}}(\operatorname{pt})\simeq\mathbb{C}[\mathfrak{t}]$ is given by the projection map $\mathcal{Z}\to\mathcal{S}\simeq\mathfrak{t}$. The grading on $H^{*}_{\mathbb{C}^{\times}}$ is determined by the weights of the $\mathbb{C}^{\times}$ representation on $\mathbb{C}[\mathcal{Z}]$. It comes from the action of $\mathbb{C}^{\times}$ on $\mathcal{Z}\subset\mathcal{S}\times X$ given by
@@
In [15] Hausel and the author extend the result to a much wider class of algebraic groups, called principally paired. The class includes all connected reductive groups and their parabolic subgroups. The affine plane $\mathcal{S}$ in general has to be replaced with a Kostant section inside the Lie algebra $\mathfrak{g}$ of the acting group $\mathrm{G}$. The torus action is analogous, and one gets the affine zero scheme $\mathcal{Z}$ over $\operatorname{Spec}H^{*}_{\mathrm{G}}(\operatorname{pt})$, and hence $\mathbb{C}[\mathcal{Z}]$ is a graded algebra over $H^{*}_{\mathrm{G}}(\operatorname{pt})$, proved to be canonically isomorphic to $H^{*}_{\mathrm{G}}(X)$. As before, the assumption is that $X$ is smooth projective, and the regular nilpotent has a unique zero. A version for singular varieties is also proved based on an analogous result of [5]. The assumptions require the singular variety to be embedded in a smooth variety with regular action, and the restriction on cohomology rings has to be surjective.
@@
First, in Section 2.2, we show that the regularity assumption can be relaxed. We only ask for the zero scheme $\mathcal{Z}$ to be of the same dimension as the base $\mathcal{S}$, i.e equal to the rank of $\mathrm{G}$. We prove in Theorem 2.7 that the ring of functions on $\mathcal{Z}$ is still canonically isomorphic to the equivariant cohomology ring, as a graded algebra over $H^{*}_{\mathrm{G}}(\operatorname{pt})$. However, the cost of this generalisation is losing the affineness of $\mathcal{Z}$, which only holds true if the action is regular. It then poses computational difficulties to potential calculations on cohomology rings. The examples are however plentiful, as we show that every spherical variety satisfies the conditions.
@@
In the same spirit, the results on equivariant cohomology in [15] were shown to hold true both for regular varieties and for GKM spaces, but without a clear connection between the two cases. However, we hope that the two results have a common generalisation into one framework, both in cohomology and in K-theory case.
@@ 1.1 Notation
All the algebraic schemes in this paper, and in particular all the algebraic groups, are defined over the complex numbers. Therefore for any scheme $X$, by $\mathbb{C}[X]=\Gamma(X,\mathcal{O}_{X})$ we denote the ring of global functions on $X$.
@@
For a linear algebraic group $\mathrm{G}$ with the Lie algebra $\mathfrak{g}$, we call an element $v\in\mathfrak{g}$ regular, if its centraliser $C_{\mathfrak{g}}(v)$ in $\mathfrak{g}$ has the dimension equal to the dimension of a maximal torus $\mathrm{T}$ in $\mathrm{G}$. If a regular element exists, the Lie algebra $\mathfrak{t}$ of $\mathrm{T}$ is a Cartan subalgebra, and its dimension is the rank of the group. The adjoint representation of $\mathrm{G}$ restricted to $\mathrm{T}$ splits into weight spaces. The weights appearing are elements of $\mathfrak{t}^{*}$. The regular elements of $\mathfrak{t}$ are then those that are not annihilated by any of the nonzero weights thereamong. Recall that for $\mathrm{G}$ reductive those are called the roots of the group.
@@ 2.1 A short recap on Kostant sections and equivariant cohomology
Notice that the non-triviality of $\phi$ is simply equivalent to the non-triviality of $d\phi(h)$. However, a group can still be principally paired with $e$ mapping to zero. This is the case for any torus, where any closed one-parameter subgroup defines a morphism from $\mathrm{B}_{2}$, factoring via the quotient by the unipotent part.
@@
Consider a homomorphism of $\mathfrak{b}_{2}=\operatorname{Lie}(\mathrm{B}_{2})$ to $\mathfrak{gl}_{n+1}$ given by
@@
The Kostant section $\mathcal{S}$ defined in [15, 2.6.3] for a principally paired group $\mathrm{G}$ satisfies the following conditions.
@@
Whenever an algebraic group $\mathrm{G}$ acts on a smooth variety $X$, it yields an infinitesimal action of the Lie algebra $\mathfrak{g}$ of $X$. Formally, the action map $\mathrm{G}\times X\to X$ yields the map $T\mathrm{G}\times TX\to TX$ of the total spaces of the tangent bundles. Restricting to $T_{1}\mathrm{G}\times X\simeq\mathfrak{g}\times X$ gives the Lie algebra action morphism $\mathfrak{g}\times X\to TX$.
Extending this by the identity on $\mathfrak{g}$ yields $\mathfrak{g}\times X\to\mathfrak{g}\times TX$, which is a section of the vertical tangent bundle on $\mathfrak{g}\times X$, i.e. the tangent bundle $\mathfrak{g}\times TX$ in the direction of $X$. We call this vector field $V$ the total vector field of the action [^3] The author would like to thank Jakub Löwit for suggesting this slick definition. The definition in [15] is technically more involved.. For any $y\in\mathfrak{g}$ this restricts to a vector field on $X$, which we denote by $V_{y}$. We also consider the restriction $V_{\mathcal{S}}$ of $V$ to $\mathcal{S}\times X$.
@@
It is not known to the author whether $\mathcal{Z}_{\mathrm{tot}}$ itself has to be reduced under the conditions of the theorem. We address this issue again in Sections 2.2 and 4.
@@
Now we need to see that the filtration is associated with the grading given by the weights of the $\mathbb{C}^{\times}$-action. First, for $t\in\mathbb{C}^{\times}$ let $t_{p}:\Omega^{p}_{X}\to\Omega^{p}_{X}$ be the pullback of forms along the map $t:X\to X$. We then construct an action of $\mathbb{C}^{\times}$ on the Koszul resolution via the following.
@@ 2.2 Zero schemes and equivariant cohomology for non-regular actions
The grading on $\mathbb{C}[\mathcal{Z}]$ is defined by the weight spaces of the $\mathbb{C}^{\times}$-representation thereon. It comes from the pullback of functions along the $\mathbb{C}^{\times}$-action on $\mathcal{Z}$ given by
@@
The assumption $\dim\mathcal{Z}=\dim\mathcal{S}$ implies that the generic fibres of the projection $\mathcal{Z}\to\mathcal{S}$ are finite, and hence a generic regular semisimple element has finitely many zeros. As the $\mathrm{T}$-fixed points are zeros of semisimple elements, there are only finitely many of them. If $X$ is a smooth projective variety with isolated torus fixed points, its cohomology is Tate. Hence we have proved
@@
Here the localisation $c|_{\zeta_{i}}$ is an element of $H^{*}_{\mathrm{T}}(\operatorname{pt})=\mathbb{C}[\mathfrak{t}]$ and hence a function on $\mathfrak{t}$, which we apply to $w$. We have thus defined $\rho(c)$ as a function on the closed points of $\mathcal{Z}$, depending on a choice for each point.
@@
We already know the isomorphism $H^{*}_{\mathrm{T}}(X)\cong H^{0}(\mathcal{Z}_{\mathrm{B}}^{\prime},\mathcal{O}_{\mathcal{Z}_{\mathrm{B}}^{\prime}})$ and moreover $H^{*}_{\mathrm{G}}(X)=H^{*}_{\mathrm{T}}(X)^{\mathrm{W}}$. Hence to arrive at $H^{*}_{\mathrm{G}}(X)\simeq H^{0}(\mathcal{Z},\mathcal{O}_{\mathcal{Z}})$, we only have to prove that the Weyl action on $\mathcal{Z}_{\mathrm{B}}^{\prime}$ coming from the action on $\mathfrak{t}$ is dual to the topological action on $H^{*}_{\mathrm{T}}(X)$. This follows the proof of [15, Lemma 4.2]. By convention we let $\mathrm{W}$ act on the left on $\mathfrak{t}$, and on the right on $\mathbb{C}[\mathfrak{t}]$ and $H^{*}_{\mathrm{T}}(X)$.
@@
Note also that as for a torus $\mathrm{T}$ the regular (and only) nilpotent in $\mathfrak{t}=\operatorname{Lie}(\mathrm{T})$ is 0, no nontrivial variety is regular with respect to a $\mathrm{T}$-action. Therefore Theorem 2.4 cannot be applied directly.
We have demonstrated that for a principally paired group $\mathrm{G}$ and a smooth projective $\mathrm{G}$-variety $X$, there are different levels of correspondence between the equivariant cohomology of $X$ and the ring of functions and coherent cohomology on $\mathcal{Z}$.
@@
It is not clear what should replace the question mark and work is ongoing to determine under what assumptions this result holds. In fact, Löwit [21] proves a version of the claim for affine Bott–Samelson varieties which are not regular.
@@ 3 Singular varieties
For any $c\in H^{*}_{\mathrm{T}}(X)$ the function $\rho(c)$ is defined on $(e+v,p)$ by localisation to a torus-fixed point in the same $\mathrm{G}$-orbit as $p$. Hence the values of $\rho(c)$ on the points of $\mathcal{Z}^{Y}$ are defined by localisation only to torus-fixed points in $Y$. Therefore the restriction of $\rho(-)$ to $\mathcal{Z}^{Y}$ factors through $\iota^{*}$. Notice that in this step we are using the reducedness of $\mathcal{Z}^{Y}$, as this allows us to recover a regular function from its values.
@@
A related work of Hausel [14] studies the equivariant cohomology ring, as well as equivariant intersection cohomology, of affine Schubert varieties. They are realised as rings of functions of zero schemes of commuting vector fields on a projective space. The definition of the vector fields is more involved.
@@ 4 K-theory conjecture for GKM spaces
In [15] a question was posed about extensions of Theorem 2.4 to other cohomology theories. The first non-trivial one, different from equivariant cohomology, is equivariant K-theory. Recall that the equivariant cohomology $H^{*}_{\mathrm{G}}(\operatorname{pt})$ of the point is equal to $\mathbb{C}[\mathfrak{t}]^{\mathrm{W}}=\mathbb{C}[\mathfrak{g}]^{\mathrm{G}}$. On the other hand, the equivariant algebraic K-theory $K^{0}_{\mathrm{G}}(\operatorname{pt})$ of the point is equal to the representation ring $R(\mathrm{G})$ of $\mathrm{G}$. After tensoring with $\mathbb{C}$, this becomes isomorphic to $\mathbb{C}[\mathrm{T}]^{\mathrm{W}}=\mathbb{C}[\mathrm{G}]^{\mathrm{G}}$ [9, Theorem 6.1.4].
@@
Third, one might ask again how important the regularity assumption is. This is elaborated on in Remark 4.5 below. In this section, we prove a K-theory analogue of the GKM theorem for cohomology [15, Theorem 5.16].
@@
FOr any $x\in K^{0}_{\mathrm{T}}(X)$ we can define $\rho(x)$ by its values, as $\operatorname{Fix}^{\prime}_{\mathrm{T}}(X)$ is reduced. For any $\mathrm{T}$-linearised vector bundle $\mathcal{E}$ on $X$ let
@@
Note that if $0\to\mathcal{E}\to\mathcal{F}\to\mathcal{G}\to 0$ is an exact sequence of $\mathrm{T}$-linearised vector bundles, then for any $(t,p)\in\operatorname{Fix}^{\prime}_{\mathrm{T}}(X)$ we have $\operatorname{Tr}_{\mathcal{F}_{p}}(t)=\operatorname{Tr}_{\mathcal{E}_{p}}(t)+\operatorname{Tr}_{\mathcal{G}_{p}}(t)$. Therefore the function $\rho$ is well defined on $K^{0}_{\mathrm{T}}(X)$, and by definition it always gives a regular function on $\operatorname{Fix}^{\prime}_{\mathrm{T}}(X)$.
@@
Having proved that $\tau$ is an injective map to $H\simeq K_{\mathrm{T}}(X)\otimes\mathbb{C}$, we only need to prove that $\tau\circ\rho=\operatorname{id}$. Take a class $\mathcal{E}]\in K^{0}_{\mathrm{T}}(X)$, where $\mathcal{E}$ is a $\mathrm{T}$-linearised vector bundle on $X$. Then $\tau\circ\rho([\mathcal{E}])$ is a function on $\mathrm{T}\times X^{\mathrm{T}}$ which at the point $(t,\zeta_{i})$ attains the value $\operatorname{Tr}_{\mathcal{E}_{\zeta_{i}}}(t)$. When we fix $\zeta_{i}$, this means that we restrict $\mathcal{E}$ to $\zeta_{i}$ to get a representation of $\mathrm{T}$, and then we compute the trace of $t$. But this is exactly how we define the isomorphism $K_{0}^{*}(\operatorname{pt})\cong\mathbb{C}[\mathrm{T}]$. Hence $\tau\circ\rho([\mathcal{E}])$, as an element on $H\subset\mathbb{C}[\mathrm{T}]^{s}$ is exactly the localisation of $\mathcal{E}$ to $X^{\mathrm{T}}$ – and that localisation is the given isomorphism between $K_{\mathrm{T}}(X)\otimes\mathbb{C}\cong H$. ∎
@@
In certain situations, we have constructed the ring homomorphisms $\rho_{H}:H^{*}_{\mathrm{G}}(X)\to\mathbb{C}[\mathcal{Z}_{\mathrm{tot}}^{\operatorname{red}}]$ as in Theorem 2.4 and $\rho_{K}:K^{0}_{\mathrm{G}}(X)\otimes\mathbb{C}\to\mathbb{C}[\operatorname{Fix}^{\prime}_{\mathrm{G}}(X)]$ as in Theorem 4.2. Whenever they are both well-defined, the complex analytic map $\operatorname{ch}_{\mathrm{G}}$ (4.1) is compatible with the Chern character $\operatorname{ch}:K^{0}_{\mathrm{G}}(X)\otimes\mathbb{C}\to H^{*}_{\mathrm{G}}(X)$. This means that the following diagram commutes.
@@
and of the construction of the Chern character from Chern roots.